# First-Hitting Location Laws as Boundary Observables of Drift–Diffusion Processes

### v1

@@ Abstract
We investigate first-hitting location (FHL) statistics induced by drift–diffusion processes in domains with absorbing boundaries, and examine how such boundary laws give rise to intrinsic information observables. Rather than introducing explicit encoding or decoding mechanisms, information is viewed as emerging directly from the geometry and dynamics of stochastic transport through first-passage events. Treating the FHL as the primary observable, we characterize how geometry and drift jointly shape the induced boundary measure. In diffusion-dominated regimes, the exit law exhibits scale-free, heavy-tailed spatial fluctuations along the boundary, whereas a nonzero drift component introduces an intrinsic length scale that suppresses these tails and reorganizes the exit statistics. Within a generator-based formulation, the FHL arises naturally as a boundary measure induced by an elliptic operator, allowing explicit boundary kernels to be derived analytically in canonical geometries. Planar absorbing boundaries in different ambient dimensions are examined as benchmark cases, illustrating how directed transport regularizes diffusion-driven fluctuations and induces qualitative transitions in boundary statistics. Overall, the present work provides a unified structural framework for first-hitting location laws and highlights FHL statistics as natural probes of geometry, drift, and irreversibility in stochastic transport.
@@ I Introduction
The present work focuses on planar absorbing geometries as canonical benchmarks for boundary-induced first-hitting location statistics. While the generator-based formulation and the qualitative transition between diffusion-dominated and drift-regularized regimes are not restricted to planar domains, explicit closed-form boundary kernels are derived here only for the half-space geometry. Extensions to curved boundaries, time-dependent drift fields, and fully information-theoretic performance measures such as channel capacity are beyond the scope of this study and will be addressed elsewhere. Our aim is instead to isolate the minimal structural mechanisms by which geometry and drift jointly shape boundary exit laws and their induced observables.
@@
Parabolic PDE route Generator–Green-function route
@@ II Model
Throughout the paper, $d$ denotes the spatial dimension; its precise interpretation is clear from context. We work in an ambient space $\mathbb{R}^{d}$ with $d\geq 2$. Particle trajectories evolve in a domain $\Omega\subset\mathbb{R}^{d}$ with absorbing boundary $\partial\Omega$. The boundary is therefore a (d1)(d-1)-dimensional manifold, and boundary-parallel coordinates are denoted by $\bm{r}\in\mathbb{R}^{d-1}$. A summary of notation is provided in Table II for reference.
@@ III A unified framework for determining first-hitting location laws
This section presents a unified route, summarized in Table I, for determining FHL laws induced by drift–diffusion processes in domains with absorbing boundaries. Related generator-based descriptions of boundary observables and boundary local time have been extensively studied in the context of reflected and absorbed diffusions; see, e.g., [10].
@@ IV.3 Three dimensional case: absorbing plane
IV.4 General $(d+1)$-dimensional structure
The explicit results in two and three dimensions reveal a common, dimension-independent structure of boundary exit laws in the absorbing half-space. In both cases, the ambient space has dimension $d+1$, while the induced boundary kernel lives on a $d$-dimensional absorbing interface. Guided by these expressions, we are naturally led to the conjecture that in $\mathbb{R}^{d+1}$ the boundary kernel admits the universal form
@@ IV.4 General $(d+1)$-dimensional structure
While (IV.4) is stated as a conjecture, its structure is strongly constrained by symmetry and scaling considerations. In particular, translational invariance along the absorbing boundary, rotational symmetry in the $d$ boundary-parallel directions, and the exponential tilting induced by the drift together single out the modified Bessel kernel appearing above. From this perspective, (IV.4) should be viewed as the natural extension of the exact two- and three-dimensional results to general boundary dimension, rather than as a heuristic guess.
As a sanity check, we specialize the conjectured expression (IV.4) to the case $d=1$, corresponding to a two-dimensional ambient space with a one-dimensional absorbing boundary. In this setting, the tangential displacement reduces to a scalar $r\in\mathbb{R}$, and the modified Bessel function appearing in (IV.4) becomes $K_{1}$ [1, 19] . Substituting $d=1$ and simplifying the prefactors, the general formula reduces exactly to
@@
where $\rho=\sqrt{\|\bm{r}\|^{2}+\lambda^{2}}$ denotes the slant distance. The above form coincides exactly with the boundary kernel obtained for the absorbing plane in Sec. IV.3, cf. (30). Together with the $d=1$ case, this agreement confirms that the conjectured form (IV.4) correctly reproduces the known closed-form results in both two and three dimensions.
@@ V.1 Asymptotic structure of planar exit laws
V.2 Differential entropy of the exit location
Let $\Xi$ denote the boundary exit location,
$\Xi:=\mathbf{X}_{\tau_{\Omega}}\in\partial\Omega,\quad\tau_{\Omega}:=\inf\{t>0:\mathbf{X}_{t}\in\partial\Omega\},$ (38)
under an initial condition $\mathbf{X}_{0}=\bm{x}$. Whenever the exit law $\omega^{\bm{x}}$ is absolutely continuous with respect to the surface measure on $\partial\Omega$, we write $\omega^{\bm{x}}(\mathrm{d}\bm{y})=\mathcal{K}(\bm{x},\bm{y})\mathrm{d}S_{\bm{y}}$ and define the (differential) entropy [6] of the exit location as
@@ V.2 Differential entropy of the exit location
This quantity characterizes the intrinsic spatial uncertainty of the exit event along the absorbing boundary and serves as a natural information observable induced by the boundary-hitting statistics.
In diffusion-dominated regimes, the exit law exhibits an algebraically decaying heavy tail along the boundary. In the drift-free limit, the kernel reduces to a Cauchy-type form with power-law decay, as in (34). Despite the absence of finite moments of order one or higher, such heavy-tailed distributions possess a well-defined differential entropy that depends only logarithmically on the scale parameter. The pathological feature relevant to information transmission is therefore not entropy divergence, but rather the lack of an intrinsic spatial scale and the resulting dominance of rare but arbitrarily large tangential displacements. In this regime, moment-based characterizations and Gaussian approximations break down, and information measures become highly sensitive to tail behavior.
The introduction of drift fundamentally alters this picture. As demonstrated by the explicit planar kernels in Sec. IV, a nonzero drift component directed toward the absorbing boundary induces an exponential suppression of large boundary displacements. This screening effect suggests a characteristic length scale $\ell_{u}$, as defined in (37), which governs the effective spatial extent of the exit distribution. For $\|\bm{v}\|>0$, the exit law is no longer scale-free, and the spatial uncertainty of the exit location is controlled by $\ell_{u}$. In particular, $\ell_{u}$ sets the effective spatial support over which the exit density contributes appreciably to entropy-like functionals, so that increasing drift strength progressively localizes the boundary distribution and reduces the contribution of rare, large displacements.
This regularization mechanism is already apparent at the level of the boundary kernel itself: the emergence of an intrinsic length scale suppresses long tangential excursions and converts an intrinsically scale-free exit observable into one governed by controlled, length-scale-dependent statistics.
A fully quantitative analysis of entropy variation or input–output information measures (e.g., the Shannon capacity [6]) would require specifying an input ensemble, and therefore lies beyond the scope of the present work.
@@ VI Conclusion and discussion
This work examined boundary exit laws associated with the location of first-passage events in drift–diffusion processes on extended domains, with particular emphasis on their asymptotic structure and induced information observables. Within a generator–Green-function framework, we derived explicit boundary kernels for planar absorbing geometries and characterized the resulting exit measures across diffusion-dominated and drift-regularized regimes. Viewed from this perspective, the exit location, rather than the exit time alone, emerges as a sensitive observable through which geometry and directed transport are imprinted on boundary statistics. In the absence of drift, the exit law reduces to harmonic measure and exhibits algebraic, scale-free decay along the boundary, whereas a nonzero drift component toward the boundary induces exponential screening and introduces a characteristic length scale $\ell_{u}=\sigma^{2}/\|\bm{v}\|=1/u$. This transition is manifest directly at the level of the boundary kernel and persists across spatial dimensions.
Beyond the availability of explicit formulas, the analysis highlights a structural distinction intrinsic to first-passage exit statistics. When drift is absent, exit locations are governed solely by geometry and diffusive spreading, leading to scale-free behavior dominated by rare but long tangential excursions. The introduction of drift qualitatively alters this picture by generating an intrinsic length scale that suppresses such excursions and localizes the exit distribution. In this sense, first-hitting location statistics provide a concrete boundary manifestation of irreversibility in stochastic transport, reflecting how drift reorganizes spatial fluctuations at absorbing interfaces. Information-theoretic quantities, such as the differential entropy of the exit location, serve as convenient diagnostics of this transition, reflecting changes in boundary statistics induced jointly by stochastic transport and geometry.
The planar results presented here provide benchmark solutions that capture the leading-order structure of boundary exit laws in more general geometries, with curvature and global shape effects entering as higher-order corrections. More broadly, the framework developed in this work identifies first-hitting location statistics as a natural class of boundary-induced observables through which geometric and dynamical aspects of stochastic transport can be examined within a unified setting. Extensions to curved absorbing boundaries and to time-dependent drift constitute natural directions for future work, where additional geometric and dynamical scales are expected to further enrich the structure of boundary exit laws.
@@ Appendix A Zero-drift limit and emergence of Cauchy exit laws
Throughout the main text, the exit law on $\partial\Omega$ is written as a boundary measure $\omega^{\bm{x}}(\mathrm{d}\bm{y})=\mathcal{K}(\bm{x},\bm{y})\,\mathrm{d}S_{\bm{y}}$. In the appendices we sometimes parameterize $\partial\Omega$ by local coordinates $y\in\mathbb{R}^{d-1}$ (e.g., $y=r$ for $d=2$ and $y=\bm{r}$ for $d=3$), in which case $\mathrm{d}S_{\bm{y}}=\mathrm{d}y$ and the same law is written in probability density function (PDF) form $f_{\Xi}(y)=\mathcal{K}(\bm{x},\bm{y}(y))$. More generally, for a coordinate map $y\mapsto\bm{y}(y)$, the density takes the form $f_{\Xi}(y)=\mathcal{K}(\bm{x},\bm{y}(y))\,J(y)$, where $J(y)$ denotes the surface Jacobian.

### v2

@@ Abstract
We investigate first-hitting location (FHL) statistics induced by drift–diffusion processes in domains with absorbing boundaries, and examine how such boundary laws give rise to intrinsic information observables. Rather than introducing explicit encoding or decoding mechanisms, information is viewed as emerging directly from the geometry and dynamics of stochastic transport through first-passage events. Treating the FHL as the primary observable, we characterize how geometry and drift jointly shape the induced boundary measure. In diffusion-dominated regimes, the exit law exhibits scale-free, heavy-tailed spatial fluctuations along the boundary, whereas a nonzero drift component introduces an intrinsic length scale that suppresses these tails and reorganizes the exit statistics. Within a generator-based formulation, the FHL arises naturally as a boundary measure induced by an elliptic operator, allowing exact $(d+1)$-dimensional boundary kernels to be derived analytically. Planar absorbing boundaries are examined as benchmark cases and validated via Monte Carlo simulations, illustrating how directed transport thermodynamically regularizes diffusion-driven fluctuations—quantified by a robust effective width—and induces qualitative transitions in boundary statistics. Overall, the present work provides a unified structural framework for first-hitting location laws and highlights FHL statistics as natural probes of geometry, drift, and irreversibility in stochastic transport.
@@ I Introduction
Novelty and main contributions.
Because the Green-function and exit-measure representations are closely related to classical potential-theory concepts, it is important to clarify what is newly synthesized here for the stochastic transport and statistical physics communities. The primary advance of this work lies in systematically treating the FHL itself as the primary physical observable, utilizing a rigorous stochastic framework that goes beyond macroscopic transport equations. Specifically, we highlight the following contributions:
  • Necessity of the stochastic framework: We provide a unified derivation that cleanly separates the spatial boundary measure from temporal stopping-time problems. We demonstrate that deriving the exact physical exit law requires a crucial exponential tilting conversion (mathematically equivalent to the Girsanov theorem [21]). This highlights that macroscopic deterministic PDEs alone are insufficient to naturally capture this measure, justifying our generator-based stochastic approach.
  • General $(d+1)$-dimensional closed-form kernels: We derive explicit, analytically closed-form half-space Poisson kernels for constant drift in arbitrary $(d+1)$ dimensions. This result fills a documented gap in foundational mathematical physics literature, which has primarily restricted explicit analytical evaluations to low dimensions ($d\leq 3$), see [23].
  • Thermodynamic regularization and numerical validation: We identify a drift-induced crossover between scale-free diffusion and exponentially screened transport regimes. We explicitly demonstrate how directed transport acts as a localized probe, physically compressing scale-free Cauchy heavy tails into an exponentially screened spatial footprint. These analytical phenomena are explicitly validated by particle-based Monte Carlo simulations of the underlying Langevin dynamics.
  • Robust geometric diagnostics: We introduce entropy-based diagnostics built on the exit-location distribution. By defining the “effective width” ($W_{\text{eff}}=\exp(H)$), we provide a robust, finite geometric measure of spatial dispersion that remains well-behaved even in the Cauchy limit, effectively resolving the divergence issues associated with traditional variance-based metrics.
A concise summary of these methodological shifts compared to the standard parabolic PDE route is provided in Table 2. Furthermore, to bridge the terminology gap between rigorous stochastic analysis and macroscopic physical intuition, Table 3 provides a methodological dictionary translating our stochastic formalism into standard transport concepts.
@@
The present work focuses on planar absorbing geometries as canonical benchmarks for boundary-induced first-hitting location statistics. While the generator-based formulation and the qualitative transition between diffusion-dominated and drift-regularized regimes are not restricted to planar domains, explicit closed-form boundary kernels are derived here only for the half-space geometry. Extensions to curved boundaries, time-dependent drift fields, and fully information-theoretic performance measures such as channel capacity are beyond the scope of this study and will be addressed in future work. Our aim is instead to isolate the minimal structural mechanisms by which geometry and drift jointly shape boundary exit laws and their induced observables.
Paper organization.
@@ II Model
Throughout the paper, $d$ denotes the spatial dimension; its precise interpretation is clear from context. We work in an ambient space $\mathbb{R}^{d}$ with $d\geq 2$. Particle trajectories evolve in a domain $\Omega\subset\mathbb{R}^{d}$ with absorbing boundary $\partial\Omega$. The boundary is therefore a (d1)(d-1)-dimensional manifold, and boundary-parallel coordinates are denoted by $\bm{r}\in\mathbb{R}^{d-1}$. A summary of notation is provided in Table 1 for reference.
@@
Table 1: Summary of key notation.
Symbol Description
$\Omega\subset\mathbb{R}^{d+1}$ Absorbing domain (half-space)
$\partial\Omega$ Absorbing boundary
$\bm{r}$ Tangential displacement along the boundary
$\lambda$ Normal distance to $\partial\Omega$
$d$ Spatial dimension
$\bm{x}\in\Omega$ Initial position
$\bm{y}\in\partial\Omega$ Exit location
$\mathbf{X}_{t}$ Drift–diffusion process
$\tau_{\Omega}$ First hitting time of $\partial\Omega$
$\Xi:=\mathbf{X}_{\tau_{\Omega}}$ Exit location (random variable)
$\omega^{\bm{x}}(\mathrm{d}\bm{y})$ Exit measure on $\partial\Omega$
$\mathcal{K}(\bm{x},\bm{y})$ Boundary kernel density
$\mathrm{d}S_{\bm{y}}$ Surface measure
$\bm{v}$ Constant drift vector
$\bm{u}$ Dimensionless drift
$D$ Diffusion coefficient
$\sigma$ Noise strength (SDE amplitude)
$K_{\nu}(\cdot)$ Modified Bessel function (second kind)
@@ III A unified framework for determining first-hitting location laws
Parabolic PDE route Generator–Green-function route
This section presents a unified route, summarized in Table 2, for determining FHL laws induced by drift–diffusion processes in domains with absorbing boundaries. Related generator-based descriptions of boundary observables and boundary local time have been extensively studied in the context of reflected and absorbed diffusions; see, e.g., [12]. To facilitate this discussion for the broad statistical physics community, Table 3 provides a direct mapping between the stochastic operators used in our framework and their macroscopic transport equivalents.
@@ IV.3 Three dimensional case: absorbing plane
IV.4 General $(d+1)$-dimensional exit law
The explicit results in two and three dimensions reveal a common, dimension-independent structure of boundary exit laws in the absorbing half-space. In both cases, the ambient space has dimension $d+1$, while the induced boundary kernel lives on a $d$-dimensional absorbing interface. Guided by low dimensional expressions, we are naturally led to the following general expression for the boundary kernel in $\mathbb{R}^{d+1}$:
@@ IV.4 General $(d+1)$-dimensional exit law
A concise derivation of Eq. (IV.4) based on the heat-kernel resolvent representation of the modified Helmholtz operator is provided in Appendix D. The structure of the kernel is strongly constrained by symmetry and scaling considerations. In particular, translational invariance along the absorbing boundary, rotational symmetry in the $d$ boundary-parallel directions, and the exponential tilting induced by the drift together single out the modified Bessel kernel appearing above. From this perspective, Eq. (IV.4) should be viewed as the natural extension of the exact two- and three-dimensional results to general boundary dimension.
As a sanity check, we specialize the general expression (IV.4) to the case $d=1$, corresponding to a two-dimensional ambient space with a one-dimensional absorbing boundary. In this setting, the tangential displacement reduces to a scalar $r\in\mathbb{R}$, and the modified Bessel function appearing in (IV.4) becomes $K_{1}$ [1, 22]. Substituting $d=1$ and simplifying the prefactors, the general formula reduces exactly to
@@
where $\rho=\sqrt{\|\bm{r}\|^{2}+\lambda^{2}}$ denotes the slant distance. The above form coincides exactly with the boundary kernel obtained for the absorbing plane in Sec. IV.3, cf. (30). Together with the $d=1$ case, this agreement confirms that Eq. (IV.4) correctly reproduces the known closed-form results in both two and three dimensions.
Figure 5: Numerical validation of the FHL distribution on a three-dimensional planar absorbing boundary. (a) Empirical PDF obtained from $10^{6}$ Monte Carlo trajectories of the Langevin dynamics Eq. (2). (b) Analytical boundary kernel calculated via Eq. (32). The parameters are set to $\lambda=1.0\,\mu\text{m}$, $\sigma=1.0$, and a constant drift $\mathbf{v}=[2,-3,1]\,\mu\text{m/s}$. The simulation results are in excellent agreement with the analytical prediction, accurately capturing the drift-induced anisotropy and the localized hitting profile.
IV.5 Numerical validation via Monte Carlo simulation
To validate the analytical boundary kernels derived in this section, we perform particle-based Monte Carlo simulations of the underlying Langevin dynamics Eq. (2). We consider a 3D system with an absorbing plane at $x_{3}=\lambda=1.0\,\mu\text{m}$ and a noise strength $\sigma=1.0$. A uniform drift $\mathbf{v}=[2,-3,1]\,\mu\text{m/s}$ is applied to introduce both normal and tangential transport components.
As shown in Fig. 5, the spatial distribution of the first-hitting locations obtained from $10^{6}$ independent trajectories is in excellent agreement with the analytical prediction of Eq. (30). Specifically, Fig. 5(a) displays the empirical probability density function (PDF) sampled from the simulated hitting events, while Fig. 5(b) shows the corresponding theoretical kernel. In particular, the theory accurately captures the drift-induced anisotropy and the exponential suppression of tangential excursions along the planar interface. This agreement confirms that the generator-based framework provides a robust structural description of the microscopic stochastic transport process, effectively bridging the path-wise Langevin dynamics with the macroscopic boundary exit law.
@@ V Asymptotic structure and information observables of exit laws
Figure 6: Spatial localization and effective width compression. (a) Exit density $K^{(1)}(r)$ for varying drift strengths $u$, showing the transition from scale-free Cauchy decay ($u=0$) to exponential screening. (b) Effective width $W_{\text{eff}}=\exp(H)$ as a function of drift strength $u$. The sharp decrease quantifies the drift-induced regularization of boundary fluctuations. Unlike the variance, which diverges in the Cauchy limit (red circle), $W_{\text{eff}}$ provides a robust and finite geometric measure of the spatial spread across all transport regimes.
@@ V.1 Asymptotic structure of planar exit laws
V.2 Differential entropy and effective width of the exit location
To characterize the intrinsic spatial uncertainty of boundary exit events, we treat the first-hitting location $\Xi:=\mathbf{X}_{\tau_{\Omega}}\in\partial\Omega$ as a primary observable. Under an initial condition $\mathbf{X}_{0}=\bm{x}$, the exit law $\omega^{\bm{x}}$ is absolutely continuous, allowing us to define the (differential) entropy [7] of the exit location as:
@@ V.2 Differential entropy and effective width of the exit location
Unlike temporal first-passage metrics, this information observable quantifies the spatial dispersion shaped by the interplay of directed transport and domain geometry.
In diffusion-dominated regimes, the exit law exhibits an algebraically decaying heavy tail. In the zero-drift limit ($\bm{u}=0$), the kernel reduces to a scale-free Cauchy-type form, see Eq. (34). A critical physical observation is that for such heavy-tailed distributions, traditional moment-based measures of spread (e.g., variance) either diverge or become highly sensitive to rare, long-range excursions. Differential entropy, by contrast, depends only logarithmically on the scale parameter and remains finite and well-behaved even in the Cauchy limit, offering a more robust diagnostic.
To make this measure physically concrete, we relate the entropy to a geometric length scale by defining the effective width (or effective occupied area) of the exit distribution:
$W_{\text{eff}}=\exp[H(\Xi\,|\,\bm{x})].$ (39)
In information theory, $W_{\text{eff}}$ corresponds to the volume of the typical set [7]. Physically, it characterizes the spatial footprint of the hitting pattern that contributes most to the probability mass—intuitively, the effective number of occupied bins at a fixed resolution.
The emergence of the intrinsic length scale $\ell_{u}=1/u$ fundamentally regularizes this spatial footprint. As illustrated in Fig. 6(a) for a one-dimensional absorbing boundary with a fixed initial normal distance $\lambda=1$, increasing the drift $u$ toward the boundary induces exponential tail screening, effectively “compressing” the distribution. To ensure numerical reproducibility, $W_{\text{eff}}$ is evaluated by numerically integrating the exact analytical kernel $\mathcal{K}^{(1)}(r;u,\lambda)$ over the infinite domain $(-\infty,\infty)$ using adaptive quadrature routines. As shown in Fig. 6(b), the computed $W_{\text{eff}}$ exhibits a sharp, monotonic decrease from its maximum in the scale-free Cauchy limit. This confirms that directed transport acts as a localized probe, suppressing diffusion-driven fluctuations and confining exit events to a region governed by $\ell_{u}$.
In summary, $W_{\text{eff}}$ serves as a robust geometric summary of the transition from scale-free to drift-dominated transport. While it characterizes the intrinsic spread of a single point source, a fully quantitative analysis of input–output information measures, such as the Shannon capacity [7], requires specifying an input ensemble and remains beyond the structural scope of this work.
@@ VI Conclusion and discussion
This work examined boundary exit laws associated with the location of first-passage events in drift–diffusion processes on extended domains, with particular emphasis on their asymptotic structure and induced information observables. Within a generator–Green-function framework, we derived exact $(d+1)$-dimensional boundary kernels for planar absorbing geometries and validated them via Monte Carlo simulations, characterizing the resulting exit measures across diffusion-dominated and drift-regularized regimes. Viewed from this perspective, the exit location, rather than the exit time alone, emerges as a sensitive observable through which geometry and directed transport are imprinted on boundary statistics. In the absence of drift, the exit law reduces to harmonic measure and exhibits algebraic, scale-free decay along the boundary, whereas a nonzero drift component toward the boundary induces exponential screening and introduces a characteristic length scale $\ell_{u}=\sigma^{2}/\|\bm{v}\|=1/u$. This transition is manifest directly at the level of the boundary kernel and persists across spatial dimensions.
Beyond the availability of explicit formulas, the analysis highlights a structural distinction intrinsic to first-passage exit statistics. When drift is absent, exit locations are governed solely by geometry and diffusive spreading, leading to scale-free behavior dominated by rare but long tangential excursions. The introduction of drift qualitatively alters this picture by generating an intrinsic length scale that suppresses such excursions and localizes the exit distribution. In this sense, first-hitting location statistics provide a concrete boundary manifestation of irreversibility in stochastic transport, reflecting how drift reorganizes spatial fluctuations at absorbing interfaces. Information-theoretic quantities, such as the differential entropy of the exit location and the associated effective width, serve as robust geometric diagnostics of this thermodynamic regularization. They effectively capture changes in boundary statistics induced jointly by stochastic transport and geometry, remaining well-behaved even when traditional variance-based metrics diverge.
The planar results presented here provide benchmark solutions that capture the leading-order structure of boundary exit laws in more general geometries, with curvature and global shape effects entering as higher-order corrections. More broadly, the framework developed in this work identifies first-hitting location statistics as a natural class of boundary-induced observables through which geometric and dynamical aspects of stochastic transport can be examined within a unified setting. From a practical standpoint, the numerical evaluation of these boundary kernels in arbitrary geometries connects to recent computational research [4] on modified Helmholtz operators. Extensions to curved absorbing boundaries, the implementation of boundary integral methods for general domains, and to time-dependent drift constitute natural directions for future work, where additional geometric and dynamical scales are expected to further enrich the structure of boundary exit laws.
@@ Appendix A Zero-Drift Limit and Emergence of Cauchy Exit Laws
Throughout the main text, the exit law on $\partial\Omega$ is written as a boundary measure $\omega^{\bm{x}}(\mathrm{d}\bm{y})=\mathcal{K}(\bm{x},\bm{y})\,\mathrm{d}S_{\bm{y}}$. In the appendices we sometimes parameterize $\partial\Omega$ by local coordinates $y\in\mathbb{R}^{d-1}$ (e.g., $y=r$ for $d=2$ and $y=\bm{r}$ for $d=3$), in which case $\mathrm{d}S_{\bm{y}}=\mathrm{d}y$ and the same law is written in PDF form $f_{\Xi}(y)=\mathcal{K}(\bm{x},\bm{y}(y))$. More generally, for a coordinate map $y\mapsto\bm{y}(y)$, the density takes the form $f_{\Xi}(y)=\mathcal{K}(\bm{x},\bm{y}(y))\,J(y)$, where $J(y)$ denotes the surface Jacobian.